OpenStudy (vane11):

I have 1-3 right, I can provide them if anyone wants, I'm just stuck on the 4th, my math isn't leading anywhere =/ Use the four-step process to find the slope of the tangent line to the graph of the given function at any point. f(x) = -1/8 x^2 Step 1: f(x + h) = Step 2: f(x + h) - f(x) = Step 3: (f(x + h) - f(x))/h = Step 4: f'(x) = lim(h->0): (f(x + h) - f(x))/h =

OpenStudy (tkhunny):

"four-step process"? Wow! Have you considered writing the actual expressions of the functions, rather than their symbolic expressions? Rather than "f(x)", write "-1/8 x^2" Rather than "f(x+h)", write "-1/8 (x+h)^2" Then, work a little algebra to simplify the resulting total expression.

OpenStudy (vane11):

Yes, I did the first three, but I'm having problems with my algebra on the fourth, is my question really that stupid?? I feel embarrassed now =/ 1) -1/8(x+h)^2 2)(-1/8(x+h)^2)-(-1/8x^2) 3)((-1/8(x+h)^2)-(-1/8x^2))/2

OpenStudy (tkhunny):

Where did "/2" come from? Should be "/h". Clarify. Is it \(f(x) = -\dfrac{1}{8x^{2}}\;or\;f(x) = -\dfrac{1}{8}x^{2}\)

OpenStudy (vane11):

Ah yes, typo that's an h not 2 my bad. And it's the second one

OpenStudy (tkhunny):

Okay, expand: \(-\dfrac{1}{8}\left((x-h)^{2} - x^{2}\right)\)

OpenStudy (tkhunny):

BTW - Do you have a textbook with this method in it? If so, will you tell me exactly what textbook it is?